geerky42 (geerky42):

Power P is required to lift a body a distance d at constant speed v. What power is require to lift the body at distance 2d at constant speed 3v?

OpenStudy (anonymous):

P ~ F v \[\frac{ P _{1} }{ P _{2} } = \frac{ v_{1} }{ v _{2} }\] \[P _{_{2}} = \frac{ v _{2} }{ v _{1}} \times P _{1}\] \[P _{2} = \frac{ 3v }{ v } \times P\] P2 = 3P

geerky42 (geerky42):

So distance doesn't matter? Will there be same result if the body is lifted at different distance?